# Collaborative Deterministic–Probabilistic Forecasting for Real-World Spatiotemporal Systems

### v4

@@ Abstract
Probabilistic forecasting is crucial for real-world spatiotemporal systems, such as climate, energy, and urban environments, where quantifying uncertainty is essential for informed, risk-aware decision-making. While diffusion models have shown promise in capturing complex data distributions, their application to spatiotemporal forecasting remains limited due to complex spatiotemporal dynamics and high computational demands. In this work, we propose CoST, a novel framework that Collaborates deterministic and diffusion models for SpatioTemporal forecasting. CoST formulates a mean-residual decomposition strategy: it leverages a powerful deterministic model to capture the conditional mean and a lightweight diffusion model to learn residual uncertainties This collaborative formulation simplifies learning objectives, enhances forecasting accuracy, enables uncertainty quantification, and significantly improves computational efficiency. To address spatial heterogeneity, we further design a scale-aware diffusion mechanism to guide the diffusion process. Extensive experiments across ten real-world datasets from climate, energy, communication, and urban systems show that CoST achieves 25% performance gains over state-of-the-art baselines, while significantly reducing computational cost.
@@ 1 Introduction
Real-world spatiotemporal systems underpin many critical domains, such as climate science, energy systems, communication networks, and urban environments. Accurate forecasting of the dynamics is essential for planning, resource allocation, and risk management [58, 5, 59, 51]. Existing approaches fall into two categories: deterministic and probabilistic forecasting. Deterministic methods estimate the conditional mean by minimizing MAE or MSE losses to capture spatiotemporal patterns [64, 37, 63]. In contrast, probabilistic methods aim to learn the full predictive distribution of observed data [46, 31, 62], enabling uncertainty quantification to support forecasting. This is particularly important in many domains, for example, in climate modeling and renewable energy, where assessing prediction reliability is essential for risk-aware decisions such as disaster preparedness and energy grid management [42, 54].
In this paper, we highlight the critical role of probabilistic forecasting in capturing uncertainty and improving the reliability of spatiotemporal predictions. However, it is non-trivial due to three challenges. First, these systems exhibit complex evolving dynamics, characterized by periodic trends, seasonal variations, and stochastic fluctuations [7, 62]. Second, these systems involve intricate spatiotemporal interactions and nonlinear dependencies [24, 63]. Third, real-world applications require both computationally efficient and scalable models [41, 51]. Recently, diffusion models have been widely adopted for probabilistic forecasting [57, 62, 46, 51]. Compared with existing approaches such as Generative Adversarial Networks (GANs) [20, 15] and Variational Autoencoders (VAEs) [28, 30], diffusion models offer superior capability in capturing complex data distributions while ensuring stable training [22, 53, 23]. These advantages make diffusion models a promising alternative. However, originally developed for image generation, they face inherent limitations in capturing temporal correlations in sequential data, as evidenced in video generation [66, 45, 9, 17] and time series forecasting [62, 47, 46, 50].
@@
To address this issue, recent efforts have explored incorporating temporal correlations as conditional inputs to guide the diffusion process [46, 50, 57], or injecting temporal priors into the noised data to explicitly model temporal correlations across time steps [31, 51, 62]. While these approaches improve temporal modeling, they remain constrained by the inherent limitations of the diffusion framework [46, 31, 47]. In contrast, we introduce a new perspective: rather than relying solely on diffusion models to capture the full data distribution, we propose a collaborative approach that combines a deterministic model and a diffusion model, leveraging their complementary strengths for probabilistic forecasting. Our design offers two key advantages. First, by leveraging powerful deterministic models to predict the conditional mean, it effectively captures the primary spatiotemporal patterns and benefits from advancements in established architectures. Second, instead of requiring the diffusion model to learn the full data distribution from scratch, we employ it to model the residuals, focusing its capacity on capturing uncertainty beyond the mean. This collaborative framework simplifies the learning objectives for each component and enhances both predictive accuracy and probabilistic expressiveness.
Building on this insight, we propose CoST, a novel framework that Collaborates deterministic and diffusion models for SpatioTemporal forecasting. As illustrated in Figure 1, we first leverage an advanced deterministic spatiotemporal forecasting model to estimate the conditional mean $\mathbb{E}[y|x]$, effectively capturing the regular patterns. Based on this, we model the residual distribution $p(r|x)=p((y-\mathbb{E}[y|x])|x)$ using a diffusion model, which complements the deterministic forecasting with uncertainty quantification. Since the diffusion model focuses solely on residuals, it allows us to adopt a lightweight denoising network and mitigate the computational overhead associated with multi-step diffusion processes. To address spatial heterogeneity, we quantify differences across spatial units and introduce a scale-aware diffusion mechanism. More importantly, we propose a comprehensive evaluation protocol for spatiotemporal probabilistic forecasting by incorporating metrics such as QICE and IS, rather than relying solely on traditional measures like CRPS, MAE, and RMSE. In summary, our main contributions are as follows:
@@
  • We propose CoST, a mean-residual decomposition approach that employs a deterministic model to estimate the conditional mean and a diffusion model to capture the residual distribution. We further design a scale-aware diffusion mechanism to address spatial heterogeneity.
@@ 2 Related Work
Spatiotemporal deterministic forecasting. Deterministic forecasting of spatiotemporal systems focuses on point estimation. These models are typically trained with loss functions like MSE or MAE to learn the conditional mean $\mathbb{E}[y|x]$, capturing regular patterns. Common deep learning architectures include MLP-based [49, 44, 67], CNN-based [29, 34, 64], and RNN-based [2, 33, 56, 55] models, valued for their efficiency. GNN-based methods [1, 3, 18, 25] capture spatial dependencies in graph-based data, while Transformer-based models [10, 12, 37, 61, 5] are effective at modeling complex temporal dynamics.
Spatiotemporal probabilistic forecasting. The core of probabilistic forecasting lies in modeling uncertainty, aiming to capture the full data distribution [60, 53]. This is particularly suited for modeling the stochastic nature of spatiotemporal systems. While early methods focused on Bayesian approaches, recent advances have explored generative models such as GANs [26, 48, 65], VAEs [11, 13, 68], and diffusion models [53, 8, 32]. Diffusion models, in particular, have gained traction for their ability to model complex distributions with stable training, yielding strong performance in spatiotemporal forecasting [46, 47, 50, 51].
Diffusion-based spatiotemporal probabilistic forecasting. Most diffusion-based forecasting methods formulate the task as conditional generation without explicitly modeling temporal dynamics, which hinders the generation of temporally coherent sequences [53, 16, 57, 47]. Moreover, the progressive corruption of time series during diffusion often distorts key patterns like long-term trends and periodicity, making temporal recovery difficult [62, 35]. To address this, methods such as TimeGrad [46] and TimeDiff [50] incorporate temporal embeddings as conditional inputs to enhance temporal awareness. Other approaches like NPDiff [51], TMDM [31], and Diffusion-TS [62] inject temporal priors into the diffusion process to better preserve temporal dynamics. More recently, DYffusion [47] redefines the denoising process to explicitly model temporal transitions at each diffusion step. Unlike prior methods, we avoid using diffusion to model temporal dynamics. Instead, we decouple forecasting into deterministic mean prediction and residual uncertainty estimation. The diffusion model focuses solely on the residuals, simplifying learning and allowing for a smaller denoising network, which greatly reduces the computational cost of the iterative diffusion process.
@@ 3 Preliminaries
We provide a summary of notations used in this paper in Appendix A.1 for clarity.
Spatiotemporal systems. Spatiotemporal systems underpin many domains such as climate science, energy, communication networks, and urban environments. The data recording spatiotemporal dynamics are typically represented as a tensor $\mathbf{x}\in\mathbb{R}^{T\times V\times C}$, where $T$, $V$, and $C$ denote the temporal, spatial, and feature dimensions, respectively. Depending on the spatial structure, the data can be organized as grid-structured ($V=H\times W$ ) or graph-structured (where $V$ represents the set of nodes). Given a historical context $\mathbf{x}^{co}=\mathbf{x}^{t-M+1:t}$ of length $M$, the goal is to predict future targets $\mathbf{x}^{ta}=\mathbf{x}^{t+1:t+P}$ over a horizon $P$ using a model $\mathcal{F}$.
@@
$p_{\theta}(\mathbf{x}_{0:N}^{ta}):=p(\mathbf{x}_{N}^{ta})\prod_{n=1}^{N}p_{\theta}(\mathbf{x}_{n-1}^{ta}|\mathbf{x}_{n}^{ta},\mathbf{x}_{0}^{co}),$ (2)
$p_{\theta}(\mathbf{x}_{n-1}^{ta}|\mathbf{x}_{n}^{ta}):=\mathcal{N}(\mathbf{x}_{n-1}^{ta};\mu_{\theta}(\mathbf{x}_{n}^{ta},n|\mathbf{x}_{0}^{co}),\Sigma_{\theta}(\mathbf{x}_{n}^{ta},n)),$
@@
Evaluations of probabilistic forecasting. We argue that probabilistic forecasting should be assessed from two key perspectives: Data Distribution—the predicted distribution should match the empirical distribution, and Prediction Usability—prediction intervals should achieve high coverage while remaining sharp. While metrics like CRPS, MAE and RMSE are widely used, they fail to assess: (i) the accuracy of quantile-wise coverage; (ii) whether the interval width reflects true uncertainty. To address this, we introduce Quantile Interval Coverage Error (QICE) [21] and Interval Score (IS) [19] as complementary metrics.
@@
$\text{QICE}:=\frac{1}{M_{\text{QIs}}}\sum_{m=1}^{M_{\text{QIs}}}\left|r_{m}-\frac{1}{M_{\text{QIs}}}\right|,\quad r_{m}=\frac{1}{N}\sum_{n=1}^{N}\mathds{1}_{y_{n}\geq\hat{y}_{n}^{\text{low}_{m}}}\cdot\mathds{1}_{y_{n}\leq\hat{y}_{n}^{\text{high}_{m}}},$ (4)
@@
$\text{IS}:=\frac{1}{N}\sum_{n=1}^{N}\left[(u_{n}^{{\alpha}_{CI}}-l_{n}^{{\alpha}_{CI}})+\frac{2}{{{\alpha}_{CI}}}(l_{n}^{{\alpha}_{CI}}-y_{n})\mathds{1}_{y_{n}<l_{n}^{{\alpha}_{CI}}}+\frac{2}{{{\alpha}_{CI}}}(y_{n}-u_{n}^{{\alpha}_{CI}})\mathds{1}_{y_{n}>u_{n}^{{\alpha}_{CI}}}\right],$ (5)
@@ 4.1 Theoretical Analysis of Mean-Residual Decomposition
Current diffusion-based probabilistic forecasting approaches typically employ a single diffusion model to capture the full distribution of data, incorporating both the regular spatiotemporal patterns and the random fluctuations. However, jointly modeling these components remains challenging [62]. Inspired by [38] and the Reynolds decomposition in fluid dynamics [43], we propose to decompose the spatiotemporal data $\mathbf{x}^{ta}$ as follows:
@@
where $\bm{\mu}$ is the conditional mean representing the regular patterns, and $\mathbf{r}$ is the residual representing the random variations. If the deterministic model approximates the conditional mean accurately, the expected residual becomes negligible, i.e. $\mathbb{E}[\mathbf{r}|\mathbf{x}^{co}]\approx 0$, and we can obtain that $\mathrm{var}(\mathbf{r}|\mathbf{x}^{co})=\mathrm{var}(\mathbf{x}^{ta}|\mathbf{x}^{co})$. Based on the law of total variance [4], we can express the variance of the target data and residuals as:
@@ 4.2 Mean Prediction via Deterministic Model
To capture the conditional mean $\mathbb{E}_{\theta}[\mathbf{x}^{ta}|\mathbf{x}^{co}]$, our framework leverages existing high-performance deterministic architectures, which are designed to capture complex spatiotemporal dynamics efficiently. In our main experiments, we use the STID [49] model as the backbone for mean prediction, and also validate our framework with ConvLSTM [52], STNorm [14], and iTransformer [36] to ensure its generality ( See Section 3). In the first stage of training, we pretrain the deterministic model for 50 epochs using historical conditional inputs $\mathbf{x}^{co}$ to output the mean estimate $\mathbb{E}_{\theta}[\mathbf{x}^{ta}|\mathbf{x}^{co}]$. The model is trained with the standard $\mathcal{L}_{2}$ loss:

### v5

# Collaborative Deterministic–Probabilistic Forecasting for Diverse Spatiotemporal Systems

@@ Abstract
Probabilistic forecasting is crucial for real-world spatiotemporal systems, such as climate, energy, and urban environments, where quantifying uncertainty is essential for informed, risk-aware decision-making. While diffusion models have shown promise in capturing complex data distributions, their application to spatiotemporal forecasting remains limited due to complex spatiotemporal dynamics and high computational demands. we propose CoST, a general forecasting framework that Collaborates deterministic and diffusion models for diverse SpatioTemporal systems. CoST formulates a mean-residual decomposition strategy: it leverages a powerful deterministic model to capture the conditional mean and a lightweight diffusion model to learn residual uncertainties. This collaborative formulation simplifies learning objectives, improves accuracy and efficiency, and generalizes across diverse spatiotemporal systems. To address spatial heterogeneity, we further design a scale-aware diffusion mechanism to guide the diffusion process. Extensive experiments across ten real-world datasets from climate, energy, communication, and urban systems show that CoST achieves 25% performance gains over state-of-the-art baselines, while significantly reducing computational cost. Code and datasets are available at: https://anonymous.4open.science/r/CoST_17116.
@@ 1 Introduction
Real-world spatiotemporal systems underpin many critical domains, such as climate science, energy systems, communication networks, and urban environments. Accurate forecasting of the dynamics is essential for planning, resource allocation, and risk management (Xie et al., 2020; Boussif et al., 2023; Xu et al., 2021; Sheng et al., 2025). Existing approaches fall into two categories: deterministic and probabilistic forecasting. Deterministic methods estimate the conditional mean by minimizing MAE or MSE losses to capture spatiotemporal patterns (Zhang et al., 2017; Ma et al., 2024; Yuan et al., 2024). In contrast, probabilistic methods aim to learn the full predictive distribution of observed data (Rasul et al., 2021; Li et al., 2024; Yuan & Qiao, 2024), enabling uncertainty quantification to support forecasting. This is particularly important in many domains, for example, in climate modeling and renewable energy, where assessing prediction reliability is essential for risk-aware decisions such as disaster preparedness and energy grid management (Palmer, 2012; Vargas Zeppetello et al., 2022).
In this paper, we highlight the critical role of probabilistic forecasting in capturing uncertainty and improving the reliability of spatiotemporal predictions. However, it is non-trivial due to three challenges. First, these systems exhibit complex evolving dynamics, characterized by periodic trends, seasonal variations, and stochastic fluctuations (Cao et al., 2021; Yuan & Qiao, 2024). Second, these systems involve intricate spatiotemporal interactions and nonlinear dependencies (Jiang et al., 2017; Yuan et al., 2024). Third, real-world applications require both computationally efficient and scalable models (Palmer, 2014; Sheng et al., 2025). Recently, diffusion models have been widely adopted for probabilistic forecasting (Wen et al., 2023; Yuan & Qiao, 2024; Rasul et al., 2021; Sheng et al., 2025). Compared with existing approaches such as Generative Adversarial Networks (GANs) (Goodfellow et al., 2020; Gao et al., 2022) and Variational Autoencoders (VAEs) (Kingma, 2013; Li et al., 2022), diffusion models offer superior capability in capturing complex data distributions while ensuring stable training (Ho et al., 2020; Tashiro et al., 2021; Ho et al., 2022). These advantages make diffusion models a promising alternative. However, originally developed for image generation, they face inherent limitations in capturing temporal correlations in sequential data, as evidenced in video generation (Zhang et al., 2024; Qiu et al., 2023; Chang et al., 2024; Ge et al., 2023) and time series forecasting (Yuan & Qiao, 2024; Rühling Cachay et al., 2023; Shen & Kwok, 2023).
@@
To address this issue, recent efforts have explored incorporating temporal correlations as conditional inputs to guide the diffusion process (Rasul et al., 2021; Shen & Kwok, 2023; Wen et al., 2023), or injecting temporal priors into the noised data to explicitly model temporal correlations across time steps (Li et al., 2024; Sheng et al., 2025; Yuan & Qiao, 2024). While these approaches improve temporal modeling, they remain constrained by the inherent limitations of the diffusion framework (Li et al., 2024; Rühling Cachay et al., 2023). In contrast, we introduce a new perspective: rather than relying solely on diffusion models to capture the full data distribution, we propose a collaborative approach that combines a deterministic model and a diffusion model, leveraging their complementary strengths for probabilistic forecasting. Our design offers two key advantages. First, by leveraging powerful deterministic models to predict the conditional mean, it effectively captures the primary spatiotemporal patterns and benefits from advancements in established architectures. Second, instead of requiring the diffusion model to learn the full data distribution from scratch, we employ it to model the residuals, focusing its capacity on capturing uncertainty beyond the mean. This collaborative framework simplifies the learning objectives for each component and enhances both predictive accuracy and probabilistic expressiveness.
Building on this insight, we propose CoST, a general forecasting framework that Collaborates deterministic and diffusion models for a wide range of SpatioTemporal systems. As illustrated in Figure 1, we first leverage an advanced deterministic spatiotemporal forecasting model to estimate the conditional mean $\mathbb{E}[y|x]$, effectively capturing the regular patterns. Based on this, we model the residual distribution $p(r|x)=p((y-\mathbb{E}[y|x])|x)$ using a diffusion model, which complements the deterministic forecasting with uncertainty quantification. Since the diffusion model focuses solely on residuals, it allows us to adopt a lightweight denoising network and mitigate the computational overhead associated with multi-step diffusion processes. To address spatial heterogeneity, we quantify differences across spatial units and introduce a scale-aware diffusion mechanism. More importantly, we propose a comprehensive evaluation protocol for spatiotemporal probabilistic forecasting by incorporating metrics such as QICE and IS, rather than relying solely on traditional measures like CRPS, MAE, and RMSE. In summary, our main contributions are as follows:
@@
  • We propose CoST, a mean-residual decomposition approach that employs a deterministic model to estimate the conditional mean and a diffusion model to capture the residual distribution. We further design a scale-aware diffusion mechanism to address spatial heterogeneity. CoST is broadly applicable across a wide range of critical real-world spatiotemporal systems.
@@ 2 Related Work
We have provided definitions and related work on spatiotemporal deterministic forecasting and probabilistic forecasting in Appendix B.
Diffusion-based spatiotemporal probabilistic forecasting. Most diffusion-based forecasting methods formulate the task as conditional generation without explicitly modeling temporal dynamics, which hinders the generation of temporally coherent sequences (Tashiro et al., 2021; Gao et al., 2023; Wen et al., 2023; Rühling Cachay et al., 2023). Moreover, the progressive corruption of time series during diffusion often distorts key patterns like long-term trends and periodicity, making temporal recovery difficult (Yuan & Qiao, 2024; Liu et al., 2023). To address this, methods such as TimeGrad (Rasul et al., 2021) and TimeDiff (Shen & Kwok, 2023) incorporate temporal embeddings as conditional inputs to enhance temporal awareness. Other approaches like NPDiff (Sheng et al., 2025) and Diffusion-TS (Yuan & Qiao, 2024) inject temporal priors into the diffusion process to better preserve temporal dynamics. More recently, DYffusion (Rühling Cachay et al., 2023) redefines the denoising process to explicitly model temporal transitions at each diffusion step. Unlike prior methods, we avoid using diffusion to model temporal dynamics. Instead, we decouple forecasting into deterministic mean prediction and residual uncertainty estimation, allowing the diffusion model to focus exclusively on modeling the residual distribution.
Hybrid Deterministic and Diffusion-based Forecasting Models. A promising recent trend involves hybridizing deterministic and diffusion models to achieve better results (Mardani et al., 2023; Gong et al., 2024; Li et al., 2024; Yu et al., 2024). For instance, CorrDiff (Mardani et al., 2023) refines the output of a deterministic UNet using a diffusion model for high-resolution atmospheric downscaling. Similarly, CasCast (Gong et al., 2024) adopts a cascaded pipeline where the coarse global prediction from a deterministic model guides a latent diffusion model to generate a refined full future state. In time series forecasting, TMDM (Li et al., 2024) conditions its diffusion process on the output of a transformer model to better learn the full data distribution. DiffCast (Yu et al., 2024) takes a coupled, end-to-end training approach for precipitation nowcasting, jointly optimizing a deterministic backbone with a diffusion model. In contrast, our approach, CoST, introduces a two-stage mean-residual decomposition strategy that distinctly separates mean prediction and residual distribution modeling. More importantly, while prior methods are typically tailored to a single domain, CoST is designed as a general framework that incorporates spatiotemporal considerations and residual heterogeneity, making it broadly applicable across diverse spatiotemporal systems.
@@ 3 Preliminaries
We provide a summary of notations used in this paper in Appendix C.1 for clarity.
Spatiotemporal systems. Spatiotemporal systems underpin many domains such as climate science, energy, communication networks, and urban environments.
The data recording spatiotemporal dynamics are typically represented as a tensor $\mathbf{x}\in\mathbb{R}^{T\times V\times C}$, where $T$, $V$, and $C$ denote the temporal, spatial, and feature dimensions, respectively. Depending on the spatial structure, the data can be organized as grid-structured ($V=H\times W$ ) or graph-structured (where $V$ represents the set of nodes). Given a historical context $\mathbf{x}^{co}=\mathbf{x}^{t-M+1:t}$ of length $M$, the goal is to predict future targets $\mathbf{x}^{ta}=\mathbf{x}^{t+1:t+P}$ over a horizon $P$ using a model $\mathcal{F}$.
@@
$p_{\theta}(\mathbf{x}_{0:N}^{ta})=p(\mathbf{x}_{N}^{ta})\prod_{n=1}^{N}p_{\theta}(\mathbf{x}_{n-1}^{ta}|\mathbf{x}_{n}^{ta},\mathbf{x}_{0}^{co}),$ (2)
$p_{\theta}(\mathbf{x}_{n-1}^{ta}|\mathbf{x}_{n}^{ta})=\mathcal{N}(\mathbf{x}_{n-1}^{ta};\mu_{\theta}(\mathbf{x}_{n}^{ta},n|\mathbf{x}_{0}^{co}),\Sigma_{\theta}(\mathbf{x}_{n}^{ta},n)),$
@@
Evaluations of probabilistic forecasting. We argue that probabilistic forecasting should be assessed from two key perspectives: Data Distribution—the predicted distribution should match the empirical distribution, and Prediction Usability—prediction intervals should achieve high coverage while remaining sharp. While metrics like CRPS, MAE and RMSE are widely used, they fail to assess: (i) the accuracy of quantile-wise coverage; (ii) whether the interval width reflects true uncertainty. To address this, we introduce Quantile Interval Coverage Error (QICE) (Han et al., 2022) and Interval Score (IS) (Gneiting & Raftery, 2007) as complementary metrics.
@@
$\text{QICE}=\frac{1}{M_{\text{QIs}}}\sum_{m=1}^{M_{\text{QIs}}}\left|r_{m}-\frac{1}{M_{\text{QIs}}}\right|,\quad r_{m}=\frac{1}{N}\sum_{n=1}^{N}\mathds{1}_{y_{n}\geq\hat{y}_{n}^{\text{low}_{m}}}\cdot\mathds{1}_{y_{n}\leq\hat{y}_{n}^{\text{high}_{m}}},$ (4)
@@
$\text{IS}=\frac{1}{N}\sum_{n=1}^{N}\left[(u_{n}^{{\alpha}_{CI}}-l_{n}^{{\alpha}_{CI}})+\frac{2}{{{\alpha}_{CI}}}(l_{n}^{{\alpha}_{CI}}-y_{n})\mathds{1}_{y_{n}<l_{n}^{{\alpha}_{CI}}}+\frac{2}{{{\alpha}_{CI}}}(y_{n}-u_{n}^{{\alpha}_{CI}})\mathds{1}_{y_{n}>u_{n}^{{\alpha}_{CI}}}\right],$ (5)
@@ 4.1 Theoretical Analysis of Mean-Residual Decomposition
Current diffusion-based probabilistic forecasting approaches typically employ a single diffusion model to capture the full distribution of data, incorporating both the regular spatiotemporal patterns and the random fluctuations. However, jointly modeling these components remains challenging (Yuan & Qiao, 2024). Inspired by Mardani et al. (2023) and the Reynolds decomposition in fluid dynamics (Pope, 2001), we propose to divide probabilistic forecasting into two parts: predicting the conditional mean and modeling the residual distribution. The spatiotemporal data $\mathbf{x}^{ta}$ can therefore be expressed as:
@@
where $\bm{\mu}$ is the conditional mean representing the regular patterns, and $\mathbf{r}$ is the residual representing the random variations. If the deterministic model approximates the conditional mean accurately, the expected residual becomes negligible, i.e. $\mathbb{E}[\mathbf{r}|\mathbf{x}^{co}]\approx 0$, and we can obtain that $\mathrm{var}(\mathbf{r}|\mathbf{x}^{co})=\mathrm{var}(\mathbf{x}^{ta}|\mathbf{x}^{co})$. Based on the law of total variance (Bertsekas & Tsitsiklis, 2008), we can express the variance of the target data and residuals as:
@@ 4.2 Mean Prediction via Deterministic Model
To capture the conditional mean $\mathbb{E}_{\theta}[\mathbf{x}^{ta}|\mathbf{x}^{co}]$, our framework leverages existing high-performance deterministic architectures, which are designed to capture complex spatiotemporal dynamics efficiently. In our main experiments, we use the STID (Shao et al., 2022) model as the backbone for mean prediction, and also validate our framework with ConvLSTM (Shi et al., 2015), STNorm (Deng et al., 2021), and iTransformer (Liu et al., ) to ensure its generality ( See Section 3). In the first stage of training, we pretrain the deterministic model for 50 epochs using historical conditional inputs $\mathbf{x}^{co}$ to output the mean estimate $\mathbb{E}_{\theta}[\mathbf{x}^{ta}|\mathbf{x}^{co}]$. The model is trained with the standard $\mathcal{L}_{2}$ loss: